OpenStudy (anonymous):

(2y)^3 (xy)^2 show steps

OpenStudy (anonymous):

I cant read that sorry @Luis_Rivera

OpenStudy (anonymous):

8 y^3 x^2 y^2= 8 x^2 y^5

OpenStudy (anonymous):

How did you even get to this y^3 x^2 y^2

OpenStudy (anonymous):

I want the steps sorry @Luis_Rivera

OpenStudy (anonymous):

Im so confused! Can you explain

OpenStudy (anonymous):

Thanks so much @Luis_Rivera for actually explaining it. -___- Medal taken away

Directrix (directrix):

(2y)^3 (xy)^2 = ? -------------- (2y)^3 means the product of three of whatever is inside the parentheses. So, (2y)^3 means --> 2y * 2y * 2y = 2 * y * 2 * y * 2 * y

Directrix (directrix):

2 * y * 2 * y * 2 * y is the same as 2*2*2 * y * y * y which is 8 times y to the third power. 8 y ^ 3 ---> Now for the other factor.

Directrix (directrix):

(xy)^2 means (xy) times (xy). So, (xy)^2 = x * y * x * y = x * x* *y * y. Simplified, that product is x^2 * y^2 or x^2 y^2

Directrix (directrix):

So far, (2y)^3 (xy)^2 = 8 y ^ 3 * x^2* y^2

Directrix (directrix):

8 y ^ 3 * x^2* y^2 = 8 * x^2 * y^3 * y^2 = 8 x^2 * y^5 or 8 x^2 y^5 ChanHamm

Directrix (directrix):

@ChanHamm